A balloon vendor at a street fair is using a tank of helium to fill her balloons. The tank has a volume of 102.0 L and a pressure of 102.0 atm at 23.0 °C. After a while she notices that the valve has not been closed properly. The pressure had dropped to 92.0 atm. (The tank is still at 23.0°C.) 


How many moles of gas has she lost? mol
Answer #1

Volume and Temperature are constant here.

we have P1 = 102.0 atm , V1= 102.0 L , T1= 23C = 296K

P2 = 92.0 atm , V1= 102.0L , T1 = 296K

initial moles = P1V11/RT1 = (102*102)/(0.082*296) = 428.642 moles

final moles = P2V2/RT2 = (92*102)/(0.082*296) = 386.62 moles

moles of gas she lost = 428.642 - 386.62 = 42.022 moles